Grade 12Mechanics

A ball is dropped
from the top of a cliff.The ball travels 1/3 in its last second of journey.what is the height of the cliff

Ipsita
7 Years agoGrade 12

1 Answer

chetan jangir
7 Years ago
We will solve this problem just by basic formula of uniformly accelerated motion that is S=ut+1/2at^2.So let's Start Here we can write S=1/2gt^2 as here ball is dropped (u=0) Also we know from given problem that in (t-1)seconds it's displacement is 2/3S.So 2/3S=1/2g(t-1)^2 Here only S and T are unknown So we can find it had now become a maths problem with two variables and two equations